# On the Small Jumps of Lévy Processes and the Multivariate Dickman Distribution

### v1

@@ 4 Approximating Small Jumps of Lévy processes
Here and throughout, when applied to Lévy processes $\stackrel{{\scriptstyle d}}{{\rightarrow}}$ refers to weak convergence on the space $D([0,\infty),\mathbb{R}^{d})$, which is the space of càdlàg functions from $[0,\infty)$ into $\mathbb{R}^{d}$ equipped with the Skorokhod topology. For any $\gamma\in\mathbb{R}^{d}$, we write $\gamma^{*}=\{t\gamma:t\geq 0\}$ to denote the element of $D([0,\infty),\mathbb{R}^{d})$ that maps $t$ to $t\gamma$. In particular, $0^{*}$ denotes the function that is identically zero.
@@
Now, consider the multivariate Dickman Lévy process $Y^{1}=\{Y_{t}^{1}:t\geq 0\}$ with $Y_{1}^{1}\sim\mathrm{MD}(\sigma)$. We now give conditions for the scaled truncated process $\epsilon^{-1}X^{\epsilon}$ to converge to $Y^{1}$ in distribution. For the convergence to hold, we need $M^{\epsilon}\stackrel{{\scriptstyle v}}{{\rightarrow}}D^{1}$ as $\epsilon\downarrow 0$, where $M^{\epsilon}$ is the Lévy measure of $\epsilon^{-1}X^{\epsilon}_{1}$ and $D^{1}$ is the Lévy measure of $Y^{1}_{1}$. We now give several statement that are equivalent to this. In the univariate case, a version of this result is given in Proposition 2.1 of [10]. As usual, for a set $C\in\mathfrak{B}(\mathbb{S}^{d-1})$, we write $\partial C$ to denote its boundary.
@@
Let $M^{\epsilon}$, $D^{1}$, and $\sigma$ be as described above. The following statements are equivalent:
@@
    For all $10<h<$ 1 and all $C\in\mathfrak{B}(\mathbb{S}^{d-1})$ with $\sigma(\partial C)=0$, $\nu((\epsilon h,\epsilon]C)\rightarrow\sigma(C)\log\frac{1}{h}$ as $\epsilon\downarrow 0$.
@@
We now give our main result for approximating the small jumps of Lévy processes by a Dickman Lévy process. This extends the univariate result in [10] to the multivariate case.
@@
We have $\epsilon^{-1}{X^{\epsilon}}\stackrel{{\scriptstyle d}}{{\rightarrow}}Y^{1}$ as $\epsilon\downarrow 0$ if and only if any of the equivalent conditions in Proposition 1 hold. Further, when $\sigma(\mathbb{S}^{d-1})=0$, this is equivalent to $\epsilon^{-1}X^{\epsilon}\stackrel{{\scriptstyle d}}{{\rightarrow}}0^{*}$ as $\epsilon\downarrow 0$, and when $\sigma(\mathbb{S}^{d-1})>0$, it is equivalent to
@@
We now give conditions that are easily checked in an important special case. Let $L^{1}(\mathbb{S}^{d-1},\sigma)$ be the space of Borel functions $g:\mathbb{S}^{d-1}\mapsto\mathbb{R}$ with $\int_{\mathbb{S}^{d-1}}|g(s)|\sigma(\mathrm{d}s)<\infty$. For any nonnegative $g\in L^{1}(\mathbb{S}^{d-1},\sigma)$ define the finite measure $\sigma_{g}$ by $\sigma_{g}(B)=\int_{B}g(s)\sigma(\mathrm{d}s)$ for $B\in\mathfrak{B}(\mathbb{S}^{d-1})$.
@@
Example 1.
In [25] the class of so-called generalized tempered stable distributions was introduced. This class consists of infinitely divisible distributions with no Gaussian part and a Lévy measure of the form
$\nu(B)=\int_{\mathbb{S}^{d-1}}\int_{0}^{\infty}1_{B}(rs)q(r,s)r^{-1-\alpha}\mathrm{d}r\sigma(\mathrm{d}s),\ \ \ B\in\mathfrak{B}(\mathbb{S}^{d-1}),$
where $\alpha\in(0,2)$, $\sigma$ is a finite measure on $\mathbb{S}^{d-1}$, and $q:[0,\infty)\times\mathbb{S}^{d-1}\mapsto(0,\infty)$ is a Borel function such that $q(r,\cdot)\to q_{*}(\cdot)$ in $L^{1}(\mathbb{S}^{d-1},\sigma)$ as $r\downarrow 0$ for some nonnegative $q_{*}\in L^{1}(\mathbb{S}^{d-1},\sigma)$. In the limiting case when $\alpha=0$, these immediately satisfy the assumptions of Corollary 1. This class contains all gamma distributions and many of their extensions, see Section 5 below for a detailed discussion.
Theorem 4 tells us that, under appropriate conditions, we can approximately simulate from a Lévy process by approximating the small jumps by a Dickman Lévy process and the large jumps by a compound Poisson process. To see this let $X=\{X_{t}:t\geq 0\}$ be a Lévy process with $X_{1}\sim\mathrm{ID}_{0}(\nu,\gamma)$. Note that we now allow for a nonzero drift. Set $\epsilon>0$ and let $\nu^{\epsilon}(B)=\int_{B}1_{|x|\leq\epsilon}(x)\nu(\mathrm{d}x)$ and $\tilde{\nu}^{\epsilon}(B)=\int_{B}1_{|x|>\epsilon}(x)\nu(\mathrm{d}x)$ for every $B\in\mathfrak{B}(\mathbb{R}^{d})$. It follows that
@@
where $X^{\epsilon}=\{X^{\epsilon}_{t}:t\geq 0\}$ is a Lévy process with $X_{1}\sim\mathrm{ID}_{0}(\nu^{\epsilon},0)$, $\tilde{X}^{\epsilon}=\{\tilde{X}^{\epsilon}_{t}:t\geq 0\}$ is a Lévy process with $\tilde{X}_{1}\sim\mathrm{ID}_{0}(\tilde{\nu}^{\epsilon},0)$, $X^{\epsilon}$ and $\tilde{X}^{\epsilon}$ are independent, and $\gamma^{*}=\{t\gamma:t\geq 0\}$. Here, we separated $X$ into: $X^{\epsilon}$, the process of small jumps, $\tilde{X}^{\epsilon}$, the process of large jumps, and $\gamma^{*}$, the deterministic drift process. Under appropriate conditions, when $\epsilon>0$ is small, we can approximate $X^{\epsilon}$ by $\epsilon$ times a Dickman Lévy process $Y^{1}$, which gives
@@
where $Y^{1}$ is independent of $\tilde{X}^{\epsilon}$. Following ideas in [8] about approximations of small jumps by Brownian motion, we now verify that the error in this approximation approaches 00 as $\epsilon\downarrow 0$.
@@
This suggest that, for small $\epsilon$, we can approximate $Z^{\epsilon}$ by 00 and we can approximately simulate from $X$ by independently simulating $Y^{1}$ and $\tilde{X}^{\epsilon}$ and then applying (14). Several methods for simulating $Y^{1}$ are discussed in [15]. For instance, we can simulate up to a finite time horizon $T>0$ as follows. Let $E_{1},E_{2},\dots\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Exp}(\theta)$, let $U_{1},U_{2},\dots\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}U(0,1)$, and let $\xi_{1},\xi_{2},\dots\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\sigma/\theta$, where $\theta=\sigma(\mathbb{S}^{d-1})$, be independent sequences of random variables. If $\Gamma_{i}=\sum_{k=1}^{i}E_{i}$ for $i=1,2,\dots$, then
@@
In practice, of course, the infinite sum must be truncated at some finite (possibly random) value. We now turn to the simulation of $\tilde{X}^{\epsilon}$. Toward this end, note that $\tilde{\nu}^{\epsilon}$ is a finite measure, set $\lambda^{\epsilon}=\tilde{\nu}^{\epsilon}(\mathbb{R}^{d})$, and let $\tilde{\nu}^{\epsilon}_{p}=\tilde{\nu}^{\epsilon}/\lambda^{\epsilon}$. It follows that $\tilde{\nu}^{\epsilon}_{p}$ is a probability measure. Let $W_{1},W_{2},\dots\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\tilde{\nu}^{\epsilon}_{p}$ and, independent of this, let $\{N(t):t\geq 0\}$ be a Poisson process with rate $\lambda^{\epsilon}$. It is readily checked that $\tilde{X}^{\epsilon}$ is a compound Poisson process and that
@@
see, e.g., Section 6.1 in [9] for details. Thus, the problem reduces to simulation from $\tilde{\nu}^{\epsilon}_{p}$. In the next section we discuss this in an important situation. Now, we give a version of Theorem 4 for stochastic integral processes. The proof is based on combining Theorems 1 and 4.
Theorem 5.
Let $X^{\epsilon}$ and $Y^{1}$ be as in Theorem 4 and let $F$ be any integrable kernel on ×$\mathbb{R}^{\ell\times d}$. If any of the equivalent condition in Proposition 1 hold, then
$\left\{\frac{1}{\epsilon}\int_{0}^{\infty}F(s,t)\mathrm{d}X^{\epsilon}_{s}:t\geq 0\right\}\stackrel{{\scriptstyle fdd}}{{\rightarrow}}\left\{\int_{0}^{\infty}F(s,t)\mathrm{d}Y_{s}:t\geq 0\right\}\ \ \mbox{as }\ \epsilon\downarrow 0.$
In particular, this holds for OU-processes.
@@ 5 Simulation from General Multivariate Gamma Distributions
In which Theorem 4 is applied to develop an approximate simulation method for a large class of distributions and a small simulation study is given to verify the performance of this approximation.
In this section we focus on distributions of the form $\mu=\mathrm{ID}_{0}(\nu,\gamma)$, where
@@
$p>0$, $\sigma$ is a finite Borel measure on $\mathbb{S}^{d-1}$, and $q:(0,\infty)\times\mathbb{S}^{d-1}\mapsto[0,\infty)$ is a Borel function such that for each $s\in\mathbb{S}^{d-1}$ we have $\lim_{r\downarrow 0}q(r,s)=1$ and $q(\cdot,s)$ is completely monotone. In [11] such distributions are called proper $p$-tempered 00-stable distributions, see also [12] for many properties. The complete monotonicity of $q$ along with our other assumptions implies that $q(r^{p},s)=\int_{(0,\infty)}e^{-r^{p}s}Q_{s}(\mathrm{d}v)$ for some measurable family $\{Q_{s}\}_{s\in\mathbb{S}^{d-1}}$ of probability measures on $(0,\infty)$ satisfying
@@
see Remark 1 and Corollary 1 in [11]. If we take $p=1$ and $Q_{s}=\delta_{b(s)}$ for some Borel function $b:\mathbb{S}^{d-1}\mapsto(0,\infty)$, these distributions reduce to the class of multivariate gamma distributions introduced in [22]. Motivated by this, we refer to our class as the class of generalized multivariate gamma distributions (GMGD).
@@
We can consider a slight generalization of GMGD. Specifically, we can allow for a Gaussian part and remove the requirement that $\lim_{r\downarrow 0}q(r,s)=1$ holds for each $s\in\mathbb{S}^{d-1}$. The resulting distributions are denoted $J_{0,p}$ in [18] and are called extended $p$-tempered 00-stable distributions in [12]. When $p=2$ they are called class $M$, see [17] and the references therein. When $p=1$ we get the Thorin class, which is the smallest class of distributions on $\mathbb{R}^{d}$ that is closed under convolution and weak convergence and contains the distributions of all so-called elementary gamma random variables on $\mathbb{R}^{d}$, see [3]. A related characterization for every $p>0$ is given in Theorem 4.18 of [12].
Note that $q(r^{p},s)\leq 1$, thus Remark 2 guarantees that the assumptions of Theorem 4 and Corollary 2 are satisfied and that these distributions fall into the more general class discussed in Example 1. This means that we can approximately simulate from the GMGD Lévy process $X=\{X_{t}:t\geq 0\}$ with $X_{1}\sim\mathrm{ID}_{0}(\nu,\gamma)$ and $\nu$ as in (19) by using the approximation in (14). Here, to simulate $\tilde{X}^{\epsilon}=\{\tilde{X}^{\epsilon}_{t}:t\geq 0\},$ the process of large jumps, we can either use (17) or (18). Either way, we need a way to simulate from $\tilde{\nu}^{\epsilon}_{p}$. We now develop an approach to do this.
@@
let $\lambda^{\epsilon}=\tilde{\nu}^{\epsilon}(\mathbb{R}^{d})$, and let $\tilde{\nu}^{\epsilon}_{p}=\tilde{\nu}^{\epsilon}/\lambda^{\epsilon}$ be a probability measure. Our goal is to develop an approach to simulate from $\tilde{\nu}^{\epsilon}_{p}$. We begin by defining several quantities and distributions. First, for $u>0$ define
@@
Now, define a probability measure on $\mathbb{S}^{d-1}$ by
@@
Since we are in the bivariate case, we write $X_{t}=(X_{1,t},X_{2,t})$ and $\tilde{X}^{\epsilon}_{t}=(\tilde{X}^{\epsilon}_{1,t},\tilde{X}^{\epsilon}_{2,t})$ to denote the Lévy process and the compound Poisson process of large jumps, respectively, at time $t$. It is readily checked that
@@
We can similarly calculate the means, variances, and the covariance on the components of $X_{t}$. The formulas are the same, but with $\epsilon=0$. Note that these quantities scale linearly in $t$.
Refer to caption Refer to caption
(a) (b)
Figure 1: (a) gives a simulated path of $\tilde{X}^{\epsilon}$, the compound Poisson process of large jumps. (b) gives a simulated path of $Y^{1}$, the Multivariate Dickman Lévy process used to approximate the small jumps. (c) gives the simulated path of $X=\tilde{X}^{\epsilon}+\epsilon Y^{1}$. In all plots $\epsilon=0.1$.
@@
Next, we performed a small simulation study to understand the error in our approximation. Toward this end we simulated $N=500,000$ paths of the Lévy process $X$. For each time $t$, let $m_{1}(t)$ and $s^{2}_{1}(t)$ and $m_{2}(t)$ and $s^{2}_{2}(t)$ be the sample means and sample variances for the first and second components, respectively, and let $s_{1,2}(t)$ be the sample covariance. Now, set
@@
to be the errors in our estimates. Note that we divide these by $t$ since the theoretical values scale linearly in $t$. We then combine these into one total error term given by
@@
Further, to understand the performance of Algorithm 3 in simulating $\tilde{X}^{\epsilon}$, we performed a similar simulation study. We again used $N=500,000$ paths and quantified the error analogously, but now using the formulas for the means, variances, and the covariance that are appropriate for this process.
@@
The results of these simulations are presented in Figure 2. In Figure 2(a), we can see that the error in simulating $\tilde{X}^{\epsilon}$ is small for all $\epsilon$’s considered. This not not surprising as Algorithm 3 is exact for all choices of $\epsilon$. In Figure 2(b) we see that our approximate method for simulating the process $X$ works well for small $\epsilon$. In Figure 2(c) we fix $\epsilon=0.1$ and compare the performance of our approach of taking $X\approx\tilde{X}^{\epsilon}+\epsilon Y^{1}$ against a potential approach of just removing the small jumps and taking $X\approx\tilde{X}^{\epsilon}$. We can see that the approach where we model the small jumps using a Dickman Lévy process has significantly less error.
Refer to caption Refer to caption
(a) (b)

### v2

@@ 3 Multivariate Dickman Approximations of Small Jumps
In this section we derive a limit theorem showing that a Dickman Lévy process can be used to approximate the small jumps of a large class of Lévy processes. Here and throughout, when applied to Lévy processes $\stackrel{{\scriptstyle d}}{{\rightarrow}}$ refers to weak convergence on the space $D([0,\infty),\mathbb{R}^{d})$, which is the space of càdlàg functions from $0,\infty)$ into $\mathbb{R}^{d}$ equipped with the Skorokhod topology. For any $\gamma\in\mathbb{R}^{d}$, we write $\gamma^{*}=\{t\gamma:t\geq 0\}$ to denote the element of $D([0,\infty),\mathbb{R}^{d})$ that maps $t$ to $t\gamma$. In particular, $0^{*}$ denotes the function that is identically zero.
@@
Now, consider the multivariate Dickman Lévy process $Y^{1}=\{Y_{t}^{1}:t\geq 0\}$ with $Y_{1}^{1}\sim\mathrm{MD}(\sigma)$. We will give conditions for the scaled truncated process $\epsilon^{-1}X^{\epsilon}$ to converge to $Y^{1}$ in distribution. For the convergence to hold, we need $M^{\epsilon}\stackrel{{\scriptstyle v}}{{\rightarrow}}D^{1}$ as $\epsilon\downarrow 0$. We now give several statement that are equivalent to this. In the univariate case, a version of this result is given in Proposition 2.1 of [9]. As usual, for a set $C\in\mathfrak{B}(\mathbb{S}^{d-1})$, we write $\partial C$ to denote its boundary.
@@
The following statements are equivalent:
@@
    For all $0<h<1$ and all $C\in\mathfrak{B}(\mathbb{S}^{d-1})$ with $\sigma(\partial C)=0$, $\nu((\epsilon h,\epsilon]C)\rightarrow\sigma(C)\log\frac{1}{h}$ as $\epsilon\downarrow 0$.
@@
We now give our main result for approximating the small jumps of a Lévy process by a Dickman Lévy process.
@@
We have $\epsilon^{-1}{X^{\epsilon}}\stackrel{{\scriptstyle d}}{{\rightarrow}}Y^{1}$ as $\epsilon\downarrow 0$ if and only if any of the equivalent conditions in Proposition 1 hold. Furthermore, when $\sigma=0$, this is equivalent to $\epsilon^{-1}X^{\epsilon}\stackrel{{\scriptstyle d}}{{\rightarrow}}0^{*}$ as $\epsilon\downarrow 0$, and when $\sigma\neq 0$, it is equivalent to
@@
This extends the univariate result in [9] to the multivariate case. We note that, in that paper, the result was formulated in terms of a condition that is analogous to Condition 4 with $p=1$ in our Proposition 1.
Remark 1.
Clearly, the assumption of Theorem 1 holds when $X\stackrel{{\scriptstyle d}}{{=}}Y^{1}$ and hence $\epsilon^{-1}{X^{\epsilon}}\stackrel{{\scriptstyle d}}{{\rightarrow}}X^{1}$. Moreover, in this case, the process $X^{\epsilon}$ is an $\epsilon$-multivariate Dickman Lévy process and applying (3) at the level of Lévy processes gives the stronger fact that $\epsilon^{-1}X^{\epsilon}\stackrel{{\scriptstyle d}}{{=}}X^{1}$ for each $\epsilon>0$. In this sense, multivariate Dickman Lévy processes are stable under a rescaling of their jumps. The fact that these processes arises in limit theorems for small jumps is analogous to how distributions that are stable under convolution are the ones that serve as the limits of sums of iid random variables (and, equivalently, the long and short time limits of Lévy processes), see [25], [19], [11], or [13]. For more on the idea of Dickman distributions being stable, see [15].
We now give conditions that are easily checked in an important special case. The formulation is influenced by the discussion in [29]. For a finite Borel measure $\sigma$ on $\mathbb{S}^{d-1}$, let $L^{1}(\mathbb{S}^{d-1},\sigma)$ be the space of Borel functions $g:\mathbb{S}^{d-1}\mapsto\mathbb{R}$ with $\int_{\mathbb{S}^{d-1}}|g(s)|\sigma(\mathrm{d}s)<\infty$. For any nonnegative $g\in L^{1}(\mathbb{S}^{d-1},\sigma)$ define the finite measure $\sigma_{g}$ by $\sigma_{g}(B)=\int_{B}g(s)\sigma(\mathrm{d}s)$ for $B\in\mathfrak{B}(\mathbb{S}^{d-1})$.
@@
Theorem 1 tells us that, under appropriate conditions, we can approximate the small jumps of a Lévy process using a Dickman Lévy process. We can then model its large jumps using a compound Poisson process. More formally, let $X=\{X_{t}:t\geq 0\}$ be a Lévy process with $X_{1}\sim\mathrm{ID}_{0}(\nu,\gamma)$. Note that we now allow for a nonzero drift. Set $\epsilon>0$ and let $\nu^{\epsilon}(B)=\int_{|x|\leq\epsilon}1_{B}(x)\nu(\mathrm{d}x)$ and $\tilde{\nu}^{\epsilon}(B)=\int_{|x|>\epsilon}1_{B}(x)\nu(\mathrm{d}x)$ for every $B\in\mathfrak{B}(\mathbb{R}^{d})$. It follows that
@@
where $X^{\epsilon}=\{X^{\epsilon}_{t}:t\geq 0\}$ is a Lévy process with $X^{\epsilon}_{1}\sim\mathrm{ID}_{0}(\nu^{\epsilon},0)$, $\tilde{X}^{\epsilon}=\{\tilde{X}^{\epsilon}_{t}:t\geq 0\}$ is a Lévy process with $\tilde{X}_{1}\sim\mathrm{ID}_{0}(\tilde{\nu}^{\epsilon},0)$, and $X^{\epsilon}$ and $\tilde{X}^{\epsilon}$ are independent. Here, we separated $X$ into: $X^{\epsilon}$, the process of small jumps, $\tilde{X}^{\epsilon}$, the process of large jumps, and $\gamma^{*}$, the deterministic drift process. Under the assumptions of Theorem 1, when $\epsilon>0$ is small, we can approximate $X^{\epsilon}$ by $\epsilon$ times a Dickman Lévy process $Y^{1}$, which gives
@@
where $Y^{1}$ is independent of $\tilde{X}^{\epsilon}$. Following ideas in [7] about approximations of small jumps by Brownian motion, we now verify that the error in this approximation approaches 0 as $\epsilon\downarrow 0$.
@@
This suggest that, for small $\epsilon$, we can approximate $Z^{\epsilon}$ by 0 and we can approximately simulate from $X$ by independently simulating $Y^{1}$ and $\tilde{X}^{\epsilon}$ and then applying (6). Three methods for simulating $Y^{1}$ are discussed in [17]. Two of the methods are approximate, while the third is exact (under mild assumptions) when simulating on a finite mesh. In this paper, we use the shot-noise method, which, while only approximate, allows us to simulate at every point on an interval. Specifically, we can simulate up to some finite time horizon $T>0$ as follows. Let $E_{1},E_{2},\dots\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Exp}(\theta)$, let $U_{1},U_{2},\dots\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}U(0,1)$, and let $\xi_{1},\xi_{2},\dots\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\sigma/\theta$, where $\theta=\sigma(\mathbb{S}^{d-1})$, be independent sequences of random variables. If $\Gamma_{i}=\sum_{k=1}^{i}E_{i}$ for $i=1,2,\dots$, then
@@
In practice, of course, the infinite sum must be truncated at some finite (possibly random) value, which is why this is an approximate method.
We now turn to the simulation of $\tilde{X}^{\epsilon}$. Toward this end, note that $\tilde{\nu}^{\epsilon}$ is a finite measure, set $\lambda^{\epsilon}=\tilde{\nu}^{\epsilon}(\mathbb{R}^{d})$, let $\tilde{\nu}^{\epsilon}_{p}=\tilde{\nu}^{\epsilon}/\lambda^{\epsilon}$, and note that $\tilde{\nu}^{\epsilon}_{p}$ is a probability measure. Let $W_{1},W_{2},\dots\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\tilde{\nu}^{\epsilon}_{p}$ and, independent of this, let $\{N(t):t\geq 0\}$ be a Poisson process with rate $\lambda^{\epsilon}$. It is readily checked that $\tilde{X}^{\epsilon}$ is a compound Poisson process and that
@@
see, e.g., Section 6.1 in [8] for details. Thus, the problem essentially reduces to finding efficient methods to simulate from $\tilde{\nu}^{\epsilon}_{p}$.
@@ 4 Simulation from General Multivariate Gamma Distributions
There are multiple ways to extend gamma distributions to the multivariate case. However, in some approaches the resulting distributions are not infinitely divisible, see the references in [28]. Since infinite divisibility is important for many financial applications [8], we focus on an extension that maintains this property. Specifically, [28] defines multivariate gamma distributions (MGD) as follows. A distribution $\mu$ on $\mathbb{R}^{d}$ is said to be MGD if $\mu=\mathrm{ID}_{0}(\nu,0)$ with
$\nu(B)=\int_{\mathbb{S}^{d-1}}\int_{0}^{\infty}1_{B}(rs)r^{-1}e^{-b(s)r}\mathrm{d}r\sigma(\mathrm{d}s),\ \ \ B\in\mathfrak{B}(\mathbb{R}),$ (11)
where $\sigma$ is a finite measure on $\mathbb{S}^{d-1}$ and $b:\mathbb{S}^{d-1}\mapsto(0,\infty)$ is a Borel function. One limitation of this class is that it is not closed under taking convolutions. We consider a generalization that is closed under taking convolutions and has additional flexibility.
Definition 2.
A distribution $\mu$ on $\mathbb{R}^{d}$ is said to be a general multivariate gamma distribution (GMGD) if $\mu=\mathrm{ID}_{0}(\nu,\gamma)$, where
@@
$p>0$, $\sigma$ is a finite Borel measure on $\mathbb{S}^{d-1}$, and $q:(0,\infty)\times\mathbb{S}^{d-1}\mapsto[0,\infty)$ is a Borel function such that for each $s\in\mathbb{S}^{d-1}$ we have $\lim_{r\downarrow 0}q(r,s)=1$ and $q(\cdot,s)$ is completely monotone.
In [10] such distributions are called proper $p$-tempered 0-stable distributions, see also [12] for many properties. The complete monotonicity of $q$ along with our other assumptions implies that $q(r^{p},s)=\int_{(0,\infty)}e^{-r^{p}v}Q_{s}(\mathrm{d}v)$ for some measurable family $\{Q_{s}\}_{s\in\mathbb{S}^{d-1}}$ of probability measures on $(0,\infty)$ satisfying
@@
see Remark 1 and Corollary 1 in [10]. If we take $p=1$ and $Q_{s}=\delta_{b(s)}$, the Lévy measure in (12) reduces to the one in (11). When $\gamma\in[0,\infty)^{d}$ and
$\sigma\left(\mathbb{S}^{d-1}\cap\left([0,\infty)^{d}\right)^{c}\right)=0$
the distribution $\mu$ is concentrated on $0,\infty)^{d}$ and its associated Lévy process is a multivariate subordinator, which can then be used for multivariate subordination of Brownian motion. This idea was introduced in [4], see also [6] and the references therein for a discussion of application to finance.
@@
We note that GMGD forms an important subclass of several well-known classes of distributions. Specifically, if we allow for a Gaussian part and remove the requirement that $\lim_{r\downarrow 0}q(r,s)=1$ holds for each $s\in\mathbb{S}^{d-1}$, then the resulting distributions correspond to those denoted $J_{0,p}$ in [24] and called extended $p$-tempered 0-stable distributions in [12]. When $p=2$ they are called class $M$, see [23] and the references therein. When $p=1$ we get the Thorin class, which is the smallest class of distributions on $\mathbb{R}^{d}$ that is closed under convolution and weak convergence and contains the distributions of all so-called elementary gamma random variables on $\mathbb{R}^{d}$, see [3] or [21]. A related characterization for every $p>0$ is given in Theorem 4.18 of [12].
We now show that the Dickman approximation holds for the small jumps of Lévy processes associated with GMGD. Toward this end, let $\rho(r,s)=q(r^{p},s)r^{-1}$ and $g(s)=1$. Definition 2 implies that $r\rho(r,s)=q(r^{p},s)\to 1=g(s)$ as $r\downarrow 0$. From here, the fact that $r\rho(r,s)=q(r^{p},s)\leq 1$, implies that the assumptions of Remark 2 are satisfied and thus that the result of Theorem 1 holds. This means that we can approximately simulate a GMGD Lévy process by using the approximation in (6). To simulate the process of large jumps, we can either use (9) or (10). Either way, we need a way to simulate from $\tilde{\nu}^{\epsilon}_{p}$. We now develop an approach to do this.
@@
let $\lambda^{\epsilon}=\tilde{\nu}^{\epsilon}(\mathbb{R}^{d})$, and let $\tilde{\nu}^{\epsilon}_{p}=\tilde{\nu}^{\epsilon}/\lambda^{\epsilon}$ be a probability measure. For $u>0$ define
@@
Next, define a probability measure on $\mathbb{S}^{d-1}$ by
@@
Since we are in the bivariate case, we write $X_{t}=(X_{1,t},X_{2,t})$ and $\tilde{X}^{\epsilon}_{t}=(\tilde{X}^{\epsilon}_{1,t},\tilde{X}^{\epsilon}_{2,t})$ to denote the GMGD Lévy process and the compound Poisson process of large jumps, respectively, at time $t$. It is readily checked that
@@
We can similarly calculate the means, variances, and the covariance on the components of $X_{t}$. The formulas are the same, but with $\epsilon=0$. Note that all of these quantities scale linearly in $t$.
Figure 1: (a) gives a simulated path of $\tilde{X}^{\epsilon}$, the compound Poisson process of large jumps. (b) gives a simulated path of $Y^{1}$, the multivariate Dickman Lévy process used to approximate the small jumps. (c) gives the simulated path of $X=\tilde{X}^{\epsilon}+\epsilon Y^{1}$. In plots (a) and (c) we take $\epsilon=0.1$.
@@
Next, we performed a small simulation study to better understand the error in our approximation. Toward this end we simulated $N=500,000$ paths of the Lévy process $X$. For each time $t$, let $m_{1}(t)$ and $s^{2}_{1}(t)$ and $m_{2}(t)$ and $s^{2}_{2}(t)$ be the sample means and sample variances for the first and second components, respectively, and let $s_{1,2}(t)$ be the sample covariance. Now, set
@@
to be the errors in our estimates. Note that we divide by $t$ since the theoretical values scale linearly in $t$. We then combine these into one total error term given by
@@
Furthermore, to understand the performance of Algorithm 3 in simulating $\tilde{X}^{\epsilon}$, we performed a similar simulation study. We again used $N=500,000$ paths and quantified the error analogously, but now using the formulas for the means, variances, and the covariance that are appropriate for this process.
@@
The results of these simulations are presented in Figure 2. In Figure 2(a), we can see that the error in simulating $\tilde{X}^{\epsilon}$ is small for all $\epsilon$’s considered. This is not surprising as Algorithm 3 is exact for all choices of $\epsilon$. In Figure 2(b) we see that our approximate method for simulating the process $X$ works well for small $\epsilon$. We note that the difference in the error between $\epsilon=0.1$ and $\epsilon=0.01$ is small, suggesting that $\epsilon=0.1$ is a good choice for this process. In Figure 2(c) we fix $\epsilon=0.1$ and compare the performance of our approach of taking $X\approx\tilde{X}^{\epsilon}+\epsilon Y^{1}$ against a potential approach of just removing the small jumps and taking $X\approx\tilde{X}^{\epsilon}$. We can see that the approach where we model the small jumps using a Dickman Lévy process has significantly less error.